Factor the following polynomials 2x^2-2x-12 and 4x^2-12x+9
Answer by fractalier(6550)

2x^2 - 2x - 12 =
2(x^2 - x - 6) =
2(x - 3)(x + 2)
and
4x^2 - 12x + 9 =
(2x - 3)^2